OpenStudy (anonymous):

Please help anybody?What is the volume of the cone below? A. 10164 units3 B. 924 units3 C. 3388 units3 D. 5082 units3

OpenStudy (anonymous):



OpenStudy (mathmate):

Volume of cone = \( \pi r^2 h \)

OpenStudy (anonymous):

i got 31,914.96

OpenStudy (anonymous):

Btw All the answer choices next to the numbers between units have pi

OpenStudy (mathmate):

If that's the case, do NOT multiply by \( \pi \) but DO divide by three, as in the given formula.

OpenStudy (anonymous):

ohhh, wich gives me 3,388 units^3

OpenStudy (mathmate):

In reality it should be 3388\( \pi\ units^3\), as you said, they all have a \( \pi \) next to it.

OpenStudy (anonymous):

Yes I know thank you!

OpenStudy (mathmate):

yw!

OpenStudy (anonymous):

@mathmate Can you help me with another?

OpenStudy (mathmate):

Post a new question, then you get chances of more people answering. Most people love medals, so be generous with your medals!

OpenStudy (mathmate):

Sorry, I didn't mean for me, but thank you anyway. After posting the new question, you can still tag the people who enjoyed helping you before.

OpenStudy (anonymous):

I know but you did help me, so you deserved it and you're welcome. && oh, okay thanks.

OpenStudy (mathmate):

So do post your new question!